OpenStudy (anonymous):

Divide and simplify? √250x^(16) / √2x

OpenStudy (anonymous):

is the problem \[\frac{ \sqrt{250x^16} }{ \sqrt{2x} }\]

OpenStudy (anonymous):

Yea basically.

OpenStudy (anonymous):

divide 250 by 2 and x^16/x=x^15 so \[\sqrt{125x^15}=5x^7\sqrt{5x}\]